# Implicit Reasoning in Deep Time Series Forecasting

### v1

@@ Abstract
Keywords Forecasting, Time Series Foundation Models, Implicit Reasoning, Memorization
@@ 1 Introduction
Foundation models are trained on large, diverse datasets, raising a critical question for time series forecasting: do these models generalize well because they learn underlying concepts of temporal dynamics, or do they simply memorize specific patterns seen during training? If those models rely on memorization, particularly in the form of time series pattern matching, it could lead to redundancy in the stored knowledge, parameter inefficiency, and possibly limit their ability to generalize well to out-of-distribution (OOD) data. Ideally, a TSFM should be capable of implicit reasoning, allowing it not to depend solely on memorization but to infer latent temporal dynamics. Such models would be able to generalize from fewer data points, offering enhanced parameter efficiency and robustness.
@@ 2 Related work
Several foundation models, including Chronos [3], LagLlama [17], Moirai [24], MOMENT [10], TimesFM [7], and TimeGPT [9], have been developed for time series forecasting. Some studies have also examined the impact of learning with synthetic data [3, 7]. MOMENT analyzed embeddings from synthetic sinusoids to isolate components like trends, while Chronos showed strong performance in forecasting composite series but struggled with exponential trends. These works demonstrate that TSFMs can learn distinct time series functions, though it remains unclear if their success is due to large-scale training or inherent reasoning abilities.
@@ 3.1 Implicit Reasoning Tasks
Our goal is to evaluate the ability of time series models to perform implicit reasoning or knowledge manipulation on OOD time series, using several experiments inspired by language model literature [2, 1, 22, 27, 29]. We carefully design experiments to test 3 forms of implicit reasoning: composition, comparison, and inverse search. We follow the notation from [22, 29], where ‘facts’ are formatted as (subject, relation, object). For continuous data, we adapt this to (input, function, output), with defining the relationship between the function , input , and output .
@@
We assess whether models can compose time series by applying knowledge manipulations similar to mathematical operations of addition, subtraction, and multiplication as illustrated in Fig. 1(a). We first define as a set of basis functions. Given a set of input functions, , the model must infer a set of composition functions to accurately forecast the OOD target signal, . We define a composition rule based on prior work [22]:
For example, in addition composition tasks, for all , where , encompassing trend and seasonality series. Models are trained exclusively on the individual component series, and the compositions of series are withheld for final evaluations. We train models on 30 trend and seasonality series () and evaluate them on all compositions of these series (). In addition to the aforementioned composition tasks, we evaluate model performance on compositions involving sinusoidal signals with varying trends, frequencies, amplitudes, and baselines.
@@
Comparison latent reasoning refers to a model’s ability to make comparative judgments on unseen or altered data. Our comparison task involves evaluating whether models can forecast series with component values that are greater or smaller than those in series used to train the model. Given a set of input functions, , the model must infer a set of functions that scale to accurately forecast the OOD target signal, . We define comparison rules based on [22]:
We train models on sinusoidal signals with varying trends, frequencies, amplitudes, and baselines () and evaluate them on a subset of series with smaller or larger parameter values not seen during training ().
@@
We evaluate whether models can ‘disentangle’ information through the inverse search task, which requires that models trained on composite signals are able to decompose time series into individual components during inference. Given a set of composition functions , the model must infer a set of decomposed functions to accurately forecast each of the OOD target signals . We define the inverse search rule as follows:
@@ 3.2 Data
| Task | Set | Function | Parameters |
| Composition | ID | ||
| (Add., Mult.) | OOD | ||
| Composition | ID | ||
| (Sub.) | OOD | ||
| Composition | ID | ||
| OOD | |||
| Comparison | ID | ||
| OOD | , | ||
| Inverse Search | ID | ||
| OOD |
@@ 4 Results
The PatchTST model excels at forecasting the composition of time series, particularly for addition and subtraction tasks as shown in Fig 2, outperforming all other evaluated Transformer models. However, the model’s performance on these tasks is sensitive to patch length, as shown in Fig. 3(a) in Appendix 4, indicating the critical role of patch length in the model’s knowledge manipulation capabilities.
Table 2 includes Mean Absolute Error (MAE) computed across all points within the forecast horizon and averaged across series for each task for DLinear, NHITS, and Transformer models. As shown in Table 2, DLinear had the smallest forecasting error for addition and subtraction composition tasks as well as comparison tasks, while NHITS had the smallest error for multiplication composition, function composition, and inverse search tasks. Results for each task, including standard deviations, for all 15 models are shown in Tables 4, 5, and 5 in Appendix 4. These tables also include baseline results for models trained on the first 1000 samples of OOD data. Example forecasts for composition and inverse search tasks are shown in Figs. 3 and 4 in Appendix 4. MAE formulation is provided in A.5.
@@ A.4 Model Training and Parameters
We train models with the following parameters outlined in Table 3 to ensure consistent evaluation across architectures.
@@
A.5 Metrics
We use Mean Absolute Error (MAE) to evaluate model performance. Here, refers to the forecast horizon, refers to the time point at which forecasts are generated, refers to the target signal values, and refers to the model;s predicted values:

### v3

@@ 1 Introduction
Foundation models are trained on large, diverse datasets, raising a critical question for time series forecasting: do these models generalize well because they learn underlying concepts of temporal dynamics, or do they simply memorize specific patterns seen during training? If models rely on memorization, particularly in the form of time series pattern matching, it could lead to redundant knowledge storage, parameter inefficiency, and limit their ability to generalize well to out-of-distribution (OOD) data. Ideally, a TSFM should be capable of implicit reasoning, allowing it not to depend solely on memorization but to infer latent temporal dynamics. Such models would be able to generalize from fewer data points, offering enhanced parameter efficiency and robustness.
@@ 2 Related work
Several foundation models, including Chronos [3], LagLlama [20], Moirai [27], MOMENT [11], TimesFM [7], TimeGPT [10], Timer [16], and Tiny Time Mixers [8], have been developed for time series forecasting. Some studies have also examined the impact of learning with synthetic data [3, 7]. MOMENT analyzed embeddings from synthetic sinusoids to isolate components like trends, while Chronos showed strong performance in forecasting composite series but struggled with exponential trends. These works demonstrate that TSFMs can learn distinct time series functions, though it remains unclear if their success is due to large-scale training or inherent reasoning abilities.
@@ 3.1 Implicit Reasoning Tasks
We assess the implicit reasoning capabilities of models, or their ability to apply and manipulate internalized patterns in novel situations, through experiments inspired by language model research [2, 1, 25, 30, 32]. We carefully design experiments to test 3 forms of implicit reasoning: composition, comparison, and inverse search. We follow the notation from [25, 32], where ‘facts’ are formatted as (subject, relation, object). For continuous data, we adapt this to (input, function, output), with defining the relationship between the function , input , and output . We consider two sets of basis functions: and where is the set of all constant signals with linear trends and is the set of all sinusoidal signals with amplitude , frequency , and baseline shift . We also define a set of functions consisting of compositions of basis functions: . Given a set of data-generating functions and their parameters observed during training, which are considered in-distribution (ID), denoted as , we assess a model’s ability to generalize to other data-generating functions observed only at inference, which we consider OOD, denoted as .
@@
For models trained on signals generated from basis functions considered ID, we assess whether they can accurately forecast compositions of these signals that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
Models are trained exclusively on the individual component series, and the compositions of series are withheld for final evaluations as illustrated in Fig. 1(a). We train models on 30 trend and seasonality series () and evaluate them on all compositions of these series ().
@@
Comparison latent reasoning refers to a model’s ability to make comparative judgments on unseen or altered data. For models trained on data generated from functions with parameters considered ID, we evaluate whether these models can generalize to data generated by the same functions with larger or smaller parameter values that are considered OOD. Specifically, we define the datasets as follows: . Consider a specific example:
We train models on signals with one varying parameter: () and evaluate them on signals with smaller or larger parameter values not seen during training ().
@@
We evaluate whether models can ‘disentangle’ information through the inverse search task, which requires that models trained on composite signals are able to infer decompositions of functions to accurately forecast each of the OOD target signals . Specifically, we define the datasets as follows: . Consider a specific example:
@@ 3.2 Data
| Task | Dist. | Function | Parameters |
| Composition | |||
| Addition | |||
| Subtraction | |||
| Multiplication | |||
| Composition | |||
| Function | |||
| Comparison | |||
| Inverse | |||
| Search |
@@ 4 Results
PatchTST outperformed all other evaluated Transformer models on the composition task, particularly for addition and subtraction tasks as shown in Fig 2. However, the model’s performance on these tasks is sensitive to patch length, as shown in Fig. 3(a) in Appendix 4, indicating the critical role of patch length in the model’s knowledge manipulation capabilities. Table 2 includes Mean Absolute Error (MAE) computed across all points within the forecast horizon and averaged across series for each task for DLinear, NHITS, and Transformer models. As shown in Table 2, DLinear had the smallest forecasting error for addition and subtraction composition tasks as well as comparison tasks, while NHITS had the smallest error for multiplication composition, function composition, and inverse search tasks. Results for each task, including standard deviations, for all 15 models are shown in Tables 4, 5, and 5 in Appendix 4. These tables also include baseline results for models trained on the first 1000 samples of OOD data. Example forecasts for composition and inverse search tasks are shown in Figs. 3 and 4 in Appendix 4. MAE formulation is provided in A.5.
@@ 5 Discussion
Acknowledgment
This work was partially supported by the NSF (awards 2406231 and 2427948), NIH (awards R01NS124642 and R01DK131586), DARPA (HR00112420329), and the US Army (W911NF-20-D0002). We would also like to thank Kin Gutierrez Olivares for his feedback on our manuscript.
@@ A.4 Model Training and Parameters
Models were trained using Adam optimizer [12]. We train models with the following parameters outlined in Table 3 to ensure consistent evaluation across architectures. The training loss function used was Mean Squared Error (MSE), where refers to the forecast horizon, refers to the time point at which forecasts are generated, refers to the target signal values, and refers to the model’s predicted values:
The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
@@
A.5 Evaluation Metrics
We use Mean Absolute Error (MAE) to evaluate model performance. Here, refers to the forecast horizon, refers to the time point at which forecasts are generated, refers to the target signal values, and refers to the model’s predicted values:
A.6 Broader Impacts
This work aims to enhance our understanding of whether time series models generalize well by learning the underlying concepts of temporal dynamics or merely memorizing specific patterns seen during training. We find that certain models generalize effectively in systematically orchestrated out-of-distribution scenarios, suggesting underexplored reasoning capabilities beyond simple pattern memorization. Since this research is foundational and does not involve direct applications, we do not anticipate any negative societal impacts. Instead, our findings may provide valuable insights for developing more data- and computationally efficient deep learning architectures. Additionally, our results may help define the limitations of time series models, ensuring they are not applied in contexts where poor generalization performance is anticipated.
A.7 Limitations
This study has several limitations that should be acknowledged. First, our analysis relies on synthetic data, which may not fully capture the complexities and nuances present in real-world datasets. While synthetic data allows for controlled experimentation, it may limit the generalizability of our findings to broader applications. Second, we employed consistent hyperparameters across models to ensure a fair evaluation of performance; however, different combinations of hyperparameters may yield improved results. Future work will explore the impact of hyperparameter variations on model performance.
A.8 Reproducibility Statement
All models are open-source. Models trained on synthetic data can be found in the Neuralforecast library [19]. Pretrained foundation models including, LagLlama [20], Chronos [3], TimesFM [7], and MOMENT [11], are open-source, with model weights available in the Hugging Face library [26]. All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs each with 49 GiB RAM. The synthetic datasets used in our study will be released publicly with the full paper.